(EPS with Convertible Bonds) On June 1, 2012, Bluhm Company and Amanar Company merged to form Davenport Inc. A total of 800,000 shares were issued to complete the merger. The corporation reports on a calendar year basis.

On April 1, 2014, the company issued an additional 600,000 shares of stock for cash. All 1,400,000 shares were outstanding on December 31, 2014.

Davenport Inc. also issued $600,000 of 20-year, 8% convertible bonds at par on July 1, 2014. Each $1,000 bond converts to 40 shares of common at any interest date. None of the bonds have been converted to date.

Davenport Inc is preparing its annual report for the fiscal year ending December 31, 2014, and will report earnings per share figures based upon a reported after-tax net income of $1,540,000. The tax rate is 40%.

Instructions

Determine the following for 2014.

(a) The number of shares to be used for calculating:

(1) Basic earnings per share.

(2) Diluted earnings per share.

(b) The earnings figures to be used for calculating:

(1) Basic earnings per share.
Answer #1

Answer:

Here,the corporation reports on a calendar year basis.

(a) Number of shares to be used for calculating:

(1) Basic earnings per share for 2014

800,000 shares are oustanding for the entire year as they are issued on June 2012.

Weighted average number of equity shares outstanding during the period

       = (800,000 * 12/12 ) + (600,000 * 9# /12)

       = 800,000 + 450,000

       = 1,250,000 shares

# = Since the additional shares are issued on 1 April 2014.

(2) Diluted earnings per share for 2014

    = Weighted average number of equity shares outstanding during the period + potential dilutive equity shares

Here,

Weighted average number of equity shares outstanding during the period = 1,250,000 shares

Potential dilutive equity shares = 8% convertible bonds

                                                          = [($600,000 / $1,000) * 40 * 6#/12] = 12,000 shares

= 1,250,000 shares + 12,000 shares

= 1,262,000 shares

# = Since the Convertible bonds are issued on 1 July 2014.

(b) The earnings figures to be used for calculating

(1) Basic earnings per share

                       = After tax-net income = $1,540,000

(2) Diluted earnings per share                                  

   After tax-net income      = $1,540,000, tax rate is 40%

Before tax-net income ($1,540,000 * 100/60) = $2,566,667

Add: Interest expense for 6 months ($600,000 * 8% *6/12) = $24,000

Before tax-net income = $2,590,667 ($2,566,667 + $24,000)

Less: Tax @ 40%         = ($1,036,267)

After tax-net income = $1,554,400

After tax-net income is $1,554,400 for computing diluted EPS purposes.